RESIK & Diogeness NEWS
![]() |
![]() |
![]() |
![]() |
(Week 21, 2003, 19 May - 25 May)
In the display above (click to enlarge), the solar X-ray fluence seen by RESIK (black points) is plotted atop GOES red line. Only periods with good spectral measurements are indicated. The total number of good spectral measurements is given in the title line.
Operation mode: |
1 s DGI has been used.
|
|
Calibrations made: |
No calibrations made
| |
ADS Settings: |
| |
|
(ORD1) HV: Det. A - 1480 V, Det. B - 1419 V. | (ORD3) HV: Det. A - 1389 V, Det. B - 1328 V. | ||||||
Channel |
l Band |
ADS |
Channel |
l Band |
ADS | ||
#1 | #2 | 3.37 - 3.88 Å | 115 - 195 | #1 | #2 | 1.16 - 1.29 Å | 110 - 230 |
#2 | #0 | 3.82 - 4.33 Å | 75 - 170 | #2 | #0 | 1.27 - 1.44 Å | 110 - 230 |
#3 | #3 | 4.31 - 4.89 Å | 170 - 250 | #3 | #3 | 1.44 - 1.63 Å | 110 - 230 |
#4 | #1 | 4.96 - 6.09 Å | 145 - 205 | #4 | #1 | 1.65 - 2.03 Å | 140 - 230 |
|
![]() |
Data gaps due to missing/bad telemetry: ~ 0.5 h.
Note, that the link to the catalogue page is now: http://www.cbk.pan.wroc.pl/resik_catalogue.htm.
Lost Power Supply from RESIK Backup Unit
Last Friday May 23 00:29:16 2003 UT, RESIK COMP lost contact with the spectrometer. For several following orbits, we were analyzing the COMP reports on its trials to establish connections with the spectrometer. As there is a lot of redundancy in the construction of RESIK electronics (256 possible independent connection paths) and no one was found functioning, we came to the conclusion, that the reason for the problem was the broken power supply to the spectrometer. We encountered earlier (see Weekly 10, 2003) the problem with the primary power supply unit and since March 14, 2003 (see Weekly 11, 2003) the backup unit has been used. Recent problem is most probably due to some failure of this backup unit. We suspect that one of possible reasons might be the automatic switch-OFF caused by the increased temperature inside the unit. The CORONAS-F presently is within the constantly Sun-lit phase of its orbital motion (see Weekly 06, 2003) and last Thursday-Friday the illumination reached the maximum.
Figure 1. Plots of the temperature within RESIK from a number of thermometers placed inside the spectrometer. The rise of temperature is observed immediately before the power problem took place.
Figure 2. Pattern of basic voltage behaviour immediately preceding the problem described. The fluctuations seen are typical – related to the periods of HV being ON and OFF. No warning signals can bee seen in the plots.
In order to be on a save side, we ordered on Thursday the power to be taken OFF from all the group of polish instruments including RESIK and Diogeness which indeed happened around 11:00 UT. We have prepared the reactivation sequence of commands and send it to our Russian colleagues. The commands are intended to bring back the power to RESIK and on the way to check also the health of the primary power supply. We are awaiting now for the telemetry stream from COMP in order to proceed further.
We keep fingers crossed hoping the spectrometer can be reanimated.
Prepared by: Janusz Sylwester js@cbk.pan.wroc.pl and Mirek Kowalinski mk@cbk.pan.wroc.pl
The report presented has been done in "real time" and so it may contain jargon, blunders, or trivialities. We do not have also an English native speaker in our Wroclaw group!
We would be happy to discuss problems mentioned above in more details if necessary.
RESIK data are in the open public domain and can be requested from: http://surfwww.mssl.ucl.ac.uk/surf/data_request.html.
Previous RESIK_weekly notes are in the archive: http://www.cbk.pan.wroc.pl/resik_archive.htm